OpenStudy (anonymous):

the measure of two supplementary angles are in the ratio 4:5 what are the measures of the two angles?

OpenStudy (anonymous):

y > x x/y = 4/5 => 5x = 4y => y = 5x/4 y = 2x - 60 2x - 60 = 5x/4 8x - 240 = 5x 3x = 240 x = 80 degrees y = 400/4 = 100 degrees supplementary